the sum of the squares of 2 consecutive odd integers is 130. Find the integers.
Answer by mananth(16949)

let the numbers be x & x+2
x^2+(x+2)^2=130
x^2+x^2+4x+4=130
2x^2+4x-126=0
/2
x^2+2x-63=0
x^2+9x-7x-63=0
x(x+9)-7(x+9)=0
(x+9)(x-7)=0
x=-9 OR 7
the integers are -9,-7
and
7 & 9